OpenStudy (camillabrooke):

At his son's birth, a man invested $2,000 in savings at 6% for his son's college education. Approximately how long will it take for the man’s investment to double?

OpenStudy (aveline):

How often is it being compounded?

OpenStudy (camillabrooke):

This was a two-part question. The first question was At his son's birth, a man invested $2,000 in savings at 6% for his son's college education. Approximately how much, to the nearest dollar, will be available in 19 years? Answer= $6051.20 I'm not sure if that answered your question or not, I'm really dumb when it comes to stuff like this, sorry!

OpenStudy (phi):

I would assume continuous compounding and use \[P= P_0 \ e^{r t}\]

OpenStudy (phi):

if you want twice as much you would want 2 P0 \[ 2 P_0 = P_0 \ e^{0.06t} \] or after dividing both sides by P0 (the starting amount) \[ 2 = e^{0.06t} \] "take the natural log of both sides" : \[ \ln 2 = 0.06 t \] now divide both sides by 0.06 \[ \frac{\ln 2 }{0.06} = t \] you need a calculator to get a number

OpenStudy (camillabrooke):

I got approx. 11.6

OpenStudy (phi):

yes. How did you get the answer for after 19 years ?

OpenStudy (camillabrooke):

2,000(1.06)^19

OpenStudy (phi):

ok, then you are assuming compounding yearly if we use that same compounding, then the time it takes to double is 2 = (1.06)^t take the natural log of both sides: ln 2 = t * ln(1.06) and t = (ln 2) / (ln 1.06) it's a bit longer than the first answer. I would round to 12 years (unless they want a more accurate answer)

OpenStudy (camillabrooke):

Thank you so much! I really appreciate it :)